OpenStudy (anonymous):

suppose z varies directly with x and interversley with y when x=6 and y=2 and z=15 find z when x=4 and y=p????

OpenStudy (aum):

Is the answer to your previous posting wrong?

OpenStudy (anonymous):

I just did not quite understand it... I thought I did at first but if you wouldn't mind could you walk me through it?

OpenStudy (aum):

If z varies directly with x, then z = C * x where C is the proportionality constant. Do you follow this part?

OpenStudy (anonymous):

yes

OpenStudy (aum):

If z varies inversely with y, then z = D * 1 / y where D is the proportionality constant. Follow this?

OpenStudy (aum):

That is what an INVERSE means. As y increases, z decreases. As y decreases, z increases. So they are related reciprocally or inversely.

OpenStudy (anonymous):

ok

OpenStudy (aum):

If z varies directly with x AND inversely with y then z = k * x / y where k is the proportionality constant.

OpenStudy (anonymous):

alright

OpenStudy (aum):

We are given one set of x, y and z. Substitute into the equation and find k. Put x = 6, y = 2, z = 15 in z = k * x / y and find k.

OpenStudy (anonymous):

15=k(6/2) is that right?

OpenStudy (aum):

yes. simplify and solve for k.

OpenStudy (anonymous):

how would I do that?

OpenStudy (aum):

15=k(6/2) simplify right hand side: 15 = k * 3 divide both sides by 3 5 = k or k = 5.

OpenStudy (anonymous):

eureka!

OpenStudy (aum):

Put k back in the original equation: z = 5 * x / y Now put the new set of values for x and y and find z. x = 4, y = p z = 5 * 4 / p = 20/p

OpenStudy (anonymous):

so the answer would be 20/9?

OpenStudy (aum):

In your original posting does it say find z when x=4 and y=p ? or is y supposed to be some number and you made a typo?

OpenStudy (anonymous):

whoops typo! it was supposed to be 9 thank you for being so patient with me I understand it a lot more now! :)

OpenStudy (aum):

then z = 20 / 9. You are welcome.